OpenStudy (anonymous):

7x - 5y = 6 -8x + y = 12 anyone knows how to do it? (:

hero (hero):

yup

hero (hero):

Multiply the second equation by 5 then add the equations together

OpenStudy (anonymous):

can you show me the steps? :)))

hero (hero):

You can do that, or you can do this: 7x - 5y = 6 -8x + y = 12 y = (6 - 7x)/-5 y = 12 + 8x y = y -5(12 + 8x) = 6 - 7x -60 - 40x = 6 - 7x 33x = -66 x = -66/33 x = -2 y = -4

hero (hero):

Those are the steps if you set y = y

OpenStudy (anonymous):

and what would be "x"?

OpenStudy (anonymous):

got it! thank you!!!!!!!!